OpenStudy (anonymous):

Use the quadratic formula to find the x-intercepts of the function. Round to the nearest tenth, if necessary. y = -2x2 + 5x - 1

OpenStudy (campbell_st):

the formula is \[x = \frac{-b \pm \sqrt{b^2 - 4ac}}{2a}\] in your question a = -2, b = 5 and c = -1 \[x = \frac{-5 \pm \sqrt{5^2 - 4(-2)(-1)}}{2(-2)}\] just evaluate

OpenStudy (anonymous):

x = [ -5 +_ sqrt(5^2 - 4 *( -2) * (-1)] 2*(-2)

OpenStudy (anonymous):

@Luis_Rivera... i get 5.5 and -3? Is that right?

OpenStudy (radar):

No, not right.

OpenStudy (anonymous):

Im confused

OpenStudy (radar):

You should use your calculator and solve these two expressions: 1.\[1/4(5+\sqrt{17})\]2.\[1/4(5-\sqrt{17})\] I believe that is what Luis_Rivera gave you.

OpenStudy (radar):

Both answers will be positive.

OpenStudy (anonymous):

I got 2.3 and 0.2. Is that right?

OpenStudy (radar):

That sounds more better, let me see what I get on my calc' 2.28 and 0.219 Close enough I would say. Great.

OpenStudy (anonymous):

Thanks! it starting to make more sense. (:

OpenStudy (radar):

Luis_Rivera gave a very detailed step by step solution for you, it is a very good solution method.

OpenStudy (radar):

I hope you were able to see his link.

OpenStudy (anonymous):

Yes i was. BUt was still confused. :)

OpenStudy (radar):

If you have any questions regarding his solution please ask them and I will attempt to answer.

OpenStudy (anonymous):

I just didnt get the last part.. Like if i had to square 17 and +5 then divede by 4

OpenStudy (radar):

Since you obtained correct answer, I believe you've got it now. Good luck.

OpenStudy (anonymous):

Thanks! Yeah i get it now. :D